# SAV-SE: Scene-aware Audio-Visual Speech Enhancement with Selective State Space Model

### v1

@@ I Introduction
In our daily living environments, speech signals are often distorted by various environmental background noises during their propagation. Speech enhancement (SE) is a task aiming at isolating the clean speech in the presence of noise interference, resulting in improved speech intelligibility and perceptual quality [1, 2, 3, 4]. It enables natural and effective Human-Robot Interaction (HRI) and plays a crucial role in various applications, such as hearing aids, mobile communication, automatic speech recognition [5, 6, 7], speaker verification [8], and speaker tracking [9, 10, 11]. These applications underscore the importance of SE in realistic scenarios. Traditional signal processing-based SE approaches, which are derived from the assumed properties on speech and noise, are incapable of suppressing highly non-stationary noise sources [12, 13, 14]. In the past decade, with the advent of deep learning technology and increased computational resources, supervised speech enhancement solutions has achieved great success [2].
@@ II-A Audio only Speech Enhancement
Traditional SE mainly include spectral subtraction [23], Wiener filtering [24], statistical model methods [25], and Non-negative Matrix Factorization (NMF) algorithms [26]. Due to assumptions about the statistical characteristics of speech and noise, these methds fail to handle highly non-stationary noise signals. In the past decades, Deep Learning (DL) has revolutionized the field of SE, demonstrating impressive performance over traditional schemes [2, 27].
@@ II-B Audio-Visual Speech Enhancement
The AVSS method [20] employs a dual-tower CNN architecture, where video frames (focusing on the speaker’s mouth region) and audio spectrograms are encoded separately. A multi-modal mask is then generated to separate the target voice from the mixed audio, significantly improving speech enhancement performance by exploiting the spatial alignment between the visual and auditory signals. VisualVoice [45] uses the appearance of the speaker’s face as an additional visual prior to isolating the corresponding vocal qualities. By incorporating cross-modal matching and speaker consistency losses from unlabeled video data, this approach enhances speech enhancement (SE) performance, especially in scenarios where the audio signal is heavily degraded. The model benefits from visual information to maintain speaker consistency and refine the target speech signal. FlowAVSE [46] takes a novel approach by adopting a conditional flow matching algorithm and optimizing a diffusion-based U-net architecture. This method enables high-quality speech generation based on cropped face images, demonstrating improved performance in real-world SE tasks. The integration of cooperative dual-attention mechanisms and dynamic audio-visual fusion strategies further enhances the robustness and effectiveness of AVSE, making it more adaptable to a variety of noisy environments. In [47], a score-based generative model is employed which leverages audio-visual embeddings from the self-supervised AV-HuBERT model to enhance speech quality and reduce artifacts like phonetic confusions. Specifically, it integrates layer-wise features from AV-HuBERT into a noise conditional score network, demonstrating improved performance and noise robustness. LA-VOCE [48] employs a transformer-based architecture to predict mel-spectrograms from noisy audio-visual inputs, while the second stage uses a neural vocoder (HiFi-GAN) to convert these spectrograms into clear waveform audio under low SNR conditions. All the aforementioned methods rely on the effective usage of human facial information. However, the effective integration of facial information requires not only accurate facial feature extraction but also robust alignment between the visual and audio modalities under varying conditions.
@@ II-C State Space Models
State space models (SSMs) have emerged as a powerful alternative to Transformer-based architectures for modeling long-range contextual dependencies in sequential data. Unlike Transformers, which rely heavily on self-attention mechanisms to capture interactions between distant tokens, SSMs provide a more efficient framework for processing sequences by leveraging dynamic state transitions that model temporal or spatial dependencies. This approach enables them to handle longer-range dependencies with fewer computational resources and greater scalability. A recent advancement in this area is the introduction of Mamba [39], a selective state space model (SSM) designed to overcome some of the limitations of traditional SSMs. Mamba incorporates a data-dependent SSM layer, which allows it to learn more flexible and task-specific representations. This enables Mamba to serve as a powerful backbone for a variety of language models, achieving superior performance over Transformers on several large-scale, real-world datasets. In particular, Mamba demonstrates linear scalability with respect to sequence length, making it an attractive alternative for tasks that require processing long sequences efficiently.
@@ III-A Problem Formulation
where ||$\cdot$|| ⋅ | extracts the magnitude, $\operatorname{arg}(\mathbf{S}_{t,k})$ and $\operatorname{arg}(\mathbf{Y}_{t,k})$ denote the spectral phase of the clean and noisy speech signals, respectively. The estimated mask is then applied to the noisy spectrum to attain the estimate of clean waveform $\widehat{\mathbf{s}}$ via inverse STFT (iSTFT):
@@ III-B1 Preliminaries
where $\mathbf{A}\!\in\!\mathbb{R}^{N\times N}$, $\mathbf{B}\!\in\!\mathbb{R}^{N\times 1}$, and $\mathbf{C}\!\in\!\mathbb{R}^{1\times N}$ are state matrix, input matrix, and output matrix. A discretization process is applied in advance to integrate SSM into deep learning architectures, where a timescale $\mathbf{\Delta}$ is used to transform matrices $\mathbf{A},\mathbf{B}$ to their discrete counterparts $\overline{\mathbf{A}},\overline{\mathbf{B}}$. A commonly employed technique for transformation is zero-order hold (ZOH), defined as:
@@ III-B3 Visual Encoder
We employ the pre-trained contrastive audio-visual masked autoencoder (CAV-MAE) [52] to extract the visual scene embedding. Specifically, CAV-MAE employs both contrastive learning and masked data modeling to jointly learn the coordinated audio-visual representations. This enables the generation of robust audio and visual features, offering significant benefits for multi-modal tasks that require comprehension of cross-modal interaction. To ensure temporal synchronization between video and audio, we use an upsampling layer to match the video feature length to the audio one, which is crucial for their later effective fusion. The resulting visual features are formulated as:
@@ III-B4 Audio Encoder
Thanks to the employed contrastive learning strategy in CAV-MAE, which captures the temporal and semantic consistencies between audio and video signals, we adopt the pretrained audio encoder to extract the visual scenario-aware audio embedding:
@@ III-B5 ConMamba
In Figure 2, the rectangle with gradient blue illustrates the internal architecture of ConMamba used in our proposed VC-S2E. Specifically, it is composed of four layers stacked together, which takes the combined multi-modal feature representation ${\bf X}\in\mathbb{R}^{(T_{y}+T_{a}+T_{v})\times d_{\text{model}}}$ as the input to produce the PSM estimate ${\widehat{\bf M}}\in\mathbb{R}^{T\times K}$.
@@
In particular, the BiMamba layer consists of two Mamba modules, each processing the input sequence along opposing temporal axes – forward and backward, respectively. Then, the bidirectional outputs are added, enabling the model to capture complex patterns and dependencies effectively. Let $\mathbf{U}$ denote the input sequence to a BiMamba layer. Then, the forward and backward pass can be formulated as:
@@
Figure 3: The detailed diagram of the Mamba neural layer.
Mamba Layer. In Figure 3, we illustrate the detailed workflow of the Mamba layer. Given the input feature representation denoted as $\mathbf{U}$, it is transformed through two parallel paths. The right path consists of a linear fully-connected layer followed by the SiLU activation function. The left path consists of a linear layer, a depthwise 1-D convolution unit, a SiLU activation, an SSM layer, and layer normalization in that sequence. The outputs of the two paths are then merged via element-wise multiplication, followed by a linear projection layer. Formally, the workflow of the Mamba layer can be represented as:
@@ IV-A Experimental Setups
TABLE I: Comparison results of our proposed VC-S2E method and the other approaches in PESQ ($\uparrow$) and STOI ($\uparrow$) with the SNR varies from -5 dB to 15 dB (Avg.: average).
@@
Datasets For clean speech data, we employ the LibriSpeech train-clean-100 subset [53], containing 28 5392853928\,53928 539 utterances from $251$ speakers. For noise data including noisy recording and the corresponding visual scene, we employ three commonly used datasets: MUSIC [54], AVSpeech [55], and AudioSet [56]. Specifically, the MUSIC dataset contains specialized high-quality music recordings. AVSpeech involves a wide variety of interference speakers, languages, and face poses, simulating complex multi-speaker scenarios. AudioSet includes diverse environmental sound samples. These three datasets cover various aspects of audio environments and provide a comprehensive evaluation.
IV-A1 Baselines
We choose three different state-of-the-art methods to compare with our proposed VC-S2E algorithm, including i) Conformer-based MP-SENet [57] adopts a codec architecture for SE where the encoder and decoder are bridged by Conformer. ii) AViTAR [58] is an image-depicted sound generation method that transforms sounds from one space to another by altering their scene-driven acoustic signatures. We re-implement their method to fit our proposed SAV-SE task. iii) ExtBiMamba [17] is the first exploration of Bidirectional Mamba (BiMamba) in monaural SE. Specifically, it demonstrates the superiority of BiMamba as an alternative to Transformer and Conformer in SE. It should be noted that, for fair comparison, all baselines are evaluated under the same parameter settings.
@@ IV-A1 Baselines
IV-A2 Implementation Details
Noise recordings exceeding $30$ seconds are segmented into clips no longer than 30 seconds, yielding 2 50025002\,5002 500 segments from MUSIC, 3 13031303\,1303 130 from AVSpeech, and 8 56085608\,5608 560 from AudioSet. For validation experiments, 1 00010001\,0001 000 clean speech utterances, $400$, $500$, and 1 00010001\,0001 000 noise recordings were randomly selected from the aforementioned speech and noise datasets, respectively. Each speech utterance is mixed with a random noise segment at a random Signal-to-Noise Ratio (SNR) sampled from $-20$ to $20$ dB (in 1 dB increments), to create the validation set. Clean speech data for testing are taken from the Librispeech test-clean-100 corpus. $200$ clean speech utterances are randomly picked (without replacement). $200$ noise recordings were randomly selected from the three AV datasets. Each clean speech recording is degraded by a random section from the noise recording at five SNR levels, i.e., -5 dB, 0 dB, 5 dB, 10 dB, and 15 dB. This produces 1 00010001\,0001 000 noisy AV clips for evaluation. All audio recordings are resampled at a rate of 16 kHz. The STFT spectrum is calculated using a Hanning window of 32 ms (512 samples) with a hop length of 16 ms (256 samples), resulting in a $257$-point magnitude spectrum as the acoustic input to the network. In terms of time dimension, $T_{y}$ denotes the maximum time frame within a batch, while $T_{a}$ and $T_{v}$ are fixed at 128 and 49, respectively. The model dimension, $d_{model}$, is set to 256. For Mamba layers, we employ the default hyper-parameters [39]: the state dimension $d_{state}=16$, convolution dimension $d_{conv}=4$, and the expansion factor of 2.
IV-A3 Training Strategy
@@
IV-A4 Evaluation Metrics
We employ the Perceptual Evaluation of Speech Quality (PESQ[59] and the Short-Time Objective Intelligibility (STOI[60] to evaluate enhanced speech signals. The PESQ assesses the mean opinion score (MOS) for objective speech perceptual quality, with a range from $-0.5$ to $4.5$. The STOI evaluates speech intelligibility, typically in the range between 00 and $1$.
@@ IV-A4 Evaluation Metrics
IV-B Experimental Results
Table I showcases the PESQ and STOI scores of the models across five SNR levels, with the best results are highlighted in bold. The evaluation results demonstrate that our VC-S2E model consistently achieves superior performance in all three datasets. In the MUSIC dataset, the unprocessed data have an average PESQ and STOI of 1.47 and 83.65%, respectively. By applying conformer-based MP-SENet, the resulting enhanced speech has improved PESQ and STOI to 2.54 and 91.65%, respectively. The AViTAR method, which employs a Conformer-based cross-attention AV framework, achieving a PESQ of 2.73 and a STOI of 92.81%. In contrast, our VC-S2E proposal shows superior performance by incorporating the BiMamba architecture with an average PESQ of 3.02 and STOI of 94.11%. It can also be seen that by integrating the spectrum, visual, and scenario-aware audio embedding, VC-S2E exceeds ExtBiMamba with an average PESQ and STOI of 2.77 and 93.09%, respectively. It should be highlighted that the advances of our VC-S2E is evident across all SNR levels. Furthermore, the same observations can be seen on the AVSpeech and AudioSet datasets as well, where our proposed VC-S2E achieves superior performance with the PESQ and STOI of 2.76 and 92.70% for AVSpeech, 2.59 and 92.25% for AudioSet, respectively.
@@ IV-B Experimental Results
Figure 4 illustrates the PESQ and STOI of our proposed VC-S2E method and its counterpart without visual and scenario-aware embeddings where AudioSet is used as the noise dataset. In both figures, VC-S2E consistently outperforms the other across all categories. For example, the PESQ of ‘Fire’ improves from 2.50 to 2.62 while for category ‘Child speech’, the score improves from 2.58 to 2.73. Similarly, the STOI scores (Figure 4(b)) also show significant improvements. For ‘Fire’, the STOI score increases from 91.61% to 92.57% while for ‘Child speech’, it improves from 92.53% to 93.27%. These results clearly reveal that incorporating visual and audio scenario-aware embeddings leads to substantial improvements in both perceptual quality and intelligibility across each individual audio category.

### v2

@@ I Introduction
In our daily living environments, speech signals are often distorted by various environmental background noises during their propagation. Speech enhancement (SE) is a task aiming at isolating the clean speech in the presence of noise interference, resulting in improved speech intelligibility and perceptual quality [1, 2, 3, 4]. It enables natural and effective Human-Robot Interaction (HRI) and plays a crucial role in various applications, such as hearing aids, mobile communication, automatic speech recognition [5, 6, 7, 8], speaker verification [9], and speaker tracking [10, 11, 12]. These applications underscore the importance of SE in realistic scenarios. Traditional signal processing-based SE approaches, which are derived from the assumed properties on speech and noise, are incapable of suppressing highly non-stationary noise sources [13, 14, 15, 16]. In the past decade, with the advent of deep learning technology and increased computational resources, supervised speech enhancement solutions has achieved great success [2].
@@ II-A Audio only Speech Enhancement
Traditional SE mainly include Wiener filtering [25], spectral subtraction [26], statistical model methods [27, 15], and Non-negative Matrix Factorization (NMF) algorithms [28]. Due to assumptions about the statistical characteristics of speech and noise, these methds fail to handle highly non-stationary noise signals. In the past decades, Deep Learning (DL) has revolutionized the field of SE, demonstrating impressive performance over traditional schemes [2, 29].
@@
In addition to these advanced architectures, recent methods leverage explicit noise representation and adversarial training to tackle non-stationary or unseen noise. For instance, DNE[41] employs VAD to embed noise features, NIT-CycleGAN [42] injects noise-type labels, and a two-stage VAE-GAN [43] disentangles noise latent variables before adversarial refinement. Such designs significantly improve SE robustness in real-world noisy conditions.
@@ II-B Audio-Visual Speech Enhancement
The AVSS method [22] employs a dual-tower CNN architecture, where video frames (focusing on the speaker’s mouth region) and audio spectrograms are encoded separately. A multi-modal mask is then generated to separate the target voice from the mixed audio, significantly improving speech enhancement performance by exploiting the spatial alignment between the visual and auditory signals. VisualVoice [50] uses the appearance of the speaker’s face as an additional visual prior to isolating the corresponding vocal qualities. By incorporating cross-modal matching and speaker consistency losses from unlabeled video data, this approach enhances speech enhancement (SE) performance, especially in scenarios where the audio signal is heavily degraded. The model benefits from visual information to maintain speaker consistency and refine the target speech signal. FlowAVSE [51] takes a novel approach by adopting a conditional flow matching algorithm and optimizing a diffusion-based U-net architecture [52]. This method enables high-quality speech generation based on cropped face images, demonstrating improved performance in real-world SE tasks. The integration of cooperative dual-attention mechanisms and dynamic audio-visual fusion strategies further enhances the robustness and effectiveness of AVSE, making it more adaptable to a variety of noisy environments. In [53], a score-based generative model is employed which leverages audio-visual embeddings from the self-supervised AV-HuBERT model to enhance speech quality and reduce artifacts like phonetic confusions. Specifically, it integrates layer-wise features from AV-HuBERT into a noise conditional score network, demonstrating improved performance and noise robustness. LA-VOCE [54] employs a transformer-based architecture to predict mel-spectrograms from noisy audio-visual inputs, while the second stage uses a neural vocoder (HiFi-GAN) to convert these spectrograms into clear waveform audio under low SNR conditions. All the aforementioned methods rely on the effective usage of human facial information. However, the effective integration of facial information requires not only accurate facial feature extraction but also robust alignment between the visual and audio modalities under varying conditions.
@@ II-C State Space Models
State space models (SSMs) have emerged as a powerful alternative to Transformer-based architectures for modeling long-range contextual dependencies in sequential data. Unlike Transformers, which rely heavily on self-attention mechanisms to capture interactions between distant tokens, SSMs provide a more efficient framework for processing sequences by leveraging dynamic state transitions that model temporal or spatial dependencies. This approach enables them to handle longer-range dependencies with fewer computational resources and greater scalability. A recent advancement in this area is the introduction of Mamba [44], a selective state space model (SSM) designed to overcome some of the limitations of traditional SSMs. Unlike Transformers, which rely on computationally expensive self-attention mechanisms scaling quadratically with sequence length, Mamba achieves linear scalability by leveraging efficient state-space transitions. This allows it to model long-range dependencies more effectively. Additionally, the selective scan mechanism of Mamba dynamically adjusts its state transition parameters based on input, enabling it to focus on task-relevant information while filtering out noise—a feature absent in TCNs and Transformers. Compared to TCNs, which are limited to capturing local temporal patterns, Mamba integrates global and localized features through its hybrid convolution-SSM design, balancing fine-grained detail and holistic sequence understanding.
@@ III-A Problem Formulation
where ||$\cdot$|| ⋅ | extracts the magnitude, $\operatorname{arg}(\mathbf{S}_{t,k})$ and $\operatorname{arg}(\mathbf{Y}_{t,k})$ denote the spectral phase of the clean and noisy speech signals, respectively.
Unlike existing AVSE methods using facial motions as the visual cues, which may be unavailable due to occlusions, off-camera targets, or low-resolution imagery, our approach leverages an image depicting the surrounding environment to provide noise prior for speech enhancement. This process is formulated as:
$\widehat{\mathbf{M}}=\mathbf{\Phi}(\mathbf{y},\mathbf{I}|\mathbf{\Omega})$ (2)
where $\mathbf{I}$ denotes the visual cue, $\mathbf{\Phi}$ denotes our proposed method SAV-SE network (see Sec. III-B) with $\mathbf{\Omega}$ the trainable parameters.
The estimated mask is then applied to the noisy spectrum to attain the estimate of clean waveform $\widehat{\mathbf{s}}$ via inverse STFT (iSTFT):
@@ III-B1 Preliminaries
where $\mathbf{A}\!\in\!\mathbb{R}^{N\times N}$, $\mathbf{B}\!\in\!\mathbb{R}^{N\times 1}$, and $\mathbf{C}\!\in\!\mathbb{R}^{1\times N}$ are state matrix, input matrix, and output matrix, representing continuous SSM parameters. A discretization process is applied in advance to integrate SSM into deep learning architectures, where a timescale $\mathbf{\Delta}$ is used to transform matrices $\mathbf{A},\mathbf{B}$ to their discrete counterparts $\overline{\mathbf{A}},\overline{\mathbf{B}}$. A commonly employed technique for transformation is zero-order hold (ZOH), defined as:
@@ III-B3 Visual Encoder
We employ the pre-trained contrastive audio-visual masked autoencoder (CAV-MAE) [58], which integrates both contrastive learning and masked data modeling to jointly learn coordinated audio-visual representations. Specifically, the visual encoder is trained using contrastive loss, ensuring precise alignment between visual features (e.g., noisy scenes like vehicles and birds) and their corresponding audio features while distinguishing non-corresponding pairs. To ensure temporal synchronization between video and audio, we use an re-sampling layer to match the video feature length to the audio one, which is crucial for their later effective fusion. The resulting visual features are formulated as:
@@ III-B4 Audio Encoder
The pretrained CAV-MAE enhances the robustness of audio and visual features, making it particularly effective for multi-modal tasks that require a deep understanding of cross-modal interaction. To capture the temporal and semantic consistencies between audio and video signals, we adopt the pretrained audio encoder to extract the visual scenario-aware audio embedding:
@@ III-B5 ConMamba
In Figure 2, the rectangle with gradient blue depicts the internal architecture of ConMamba in our proposed VC-S2E. Specifically, it consists of four stacked layers, which take the combined multi-modal feature representation ${\bf X}\in\mathbb{R}^{(T_{y}+T_{a}+T_{v})\times d_{\text{model}}}$ as input to produce the PSM estimate ${\widehat{\bf M}}\in\mathbb{R}^{T\times K}$. In conventional state space models, the hidden state is updated sequentially. In contrast, the BiMamba layer uses a data-dependent selective scanning mechanism, where core parameters are adaptively tuned based on the visual and audio content. When a visual token indicates a noise scene (e.g., an instrument or vehicle), bidirectional scanning updates the hidden states to associate the visual noise source with corresponding audio frequency features for targeted suppression. Additionally, depthwise separable 1D convolutions (DWConv) are integrated into each layer to precisely capture local interactions between adjacent visual and audio tokens.
@@
Figure 3: The illustrations of (a) the BiMamba layer and (b) the Mamba block.
In particular, as illustrated in Figure 3, the BiMamba layer consists of two parallel Mamba blocks, each processing the input sequence along opposing temporal axes – forward and backward, respectively. Then, the bidirectional outputs are added, enabling the model to capture complex patterns and dependencies effectively. Let $\mathbf{U}$ denote the input sequence to a BiMamba layer. Then, the forward and backward pass can be formulated as:
@@
Mamba Block. In Figure 3 (b), we illustrate the detailed workflow of the Mamba block. Given the input feature representation denoted as $\mathbf{U}$, it is transformed through two parallel paths. The right path consists of a linear fully-connected layer followed by the SiLU activation function. The left path consists of a linear layer, a depthwise 1-D convolution unit, a SiLU activation, an SSM layer, and layer normalization in that sequence. The outputs of the two paths are then merged via element-wise multiplication, followed by a linear projection layer. Formally, the workflow of the Mamba block can be represented as:
@@ IV-A Experimental Setups
TABLE I: Comparison results of our proposed VC-S2E method and the other approaches in PESQ ($\uparrow$) and STOI ($\uparrow$) with the SNR varies from -5 dB to 15 dB.
@@
IV-A1 Datasets
For clean speech data, we employ the LibriSpeech train-clean-100 subset [59], containing 28 5392853928\,53928 539 utterances from $251$ speakers. For noise data including noisy recording and the corresponding visual scene, we employ three commonly used datasets: MUSIC [60], AVSpeech [61], and AudioSet [62]. Specifically, the MUSIC dataset captures structured noise from musical instruments in controlled environments, AVSpeech provides diverse multi-speaker and multi-language scenarios with varied face poses, and AudioSet covers extensive environmental sounds, including both natural and human-made noises. These characteristics collectively align with the diverse conditions where SAV-SE is designed to excel. These three datasets cover various aspects of audio environments and provide a comprehensive evaluation. The data used in training and testing are non-overlapping, with each recording featuring unseen paired visual scenes and noisy sequences.
IV-A2 Baselines
We choose seven different state-of-the-art methods to compare with our proposed VC-S2E algorithm, including i) Conformer-based MP-SENet [63] adopts a codec architecture for SE where the encoder and decoder are bridged by Conformer. ii) AViTAR [64] is an image-depicted sound generation method that transforms sounds from one space to another by altering their scene-driven acoustic signatures. We re-implement their method to fit our proposed SAV-SE task. iii) ExtBiMamba [19] is the first exploration of Bidirectional Mamba (BiMamba) in monaural SE. Specifically, it demonstrates the superiority of BiMamba as an alternative to Transformer and Conformer in SE. iv) AVDCNN [65] uses a dual-stream CNN with late fusion for multi-task audio-visual learning, enhancing speech via visual cues. v) MLDM [66] introduces a unified multi-level distortion measure for comprehensive speech enhancement. vi) ALTKD [67] uses a U-Net-based teacher-student model, where the teacher fuses ultrasound tongue images, lip videos, and audio, distilling tongue information for high-quality enhancement with only audio and lip video needed at inference. vii) DNE [41] leverages VAD to capture noise characteristics for improved SE in non-stationary and unseen noise conditions. It should be noted that, for fair comparison, all baselines are evaluated under the same parameter settings.
@@ IV-A2 Baselines
IV-A3 Implementation Details
Noise recordings exceeding $30$ seconds are segmented into clips no longer than 30 seconds, yielding 2 50025002\,5002 500 segments from MUSIC, 3 13031303\,1303 130 from AVSpeech, and 8 56085608\,5608 560 from AudioSet. For validation experiments, 1 00010001\,0001 000 clean speech utterances, $400$, $500$, and 1 00010001\,0001 000 noise recordings were randomly selected from the aforementioned speech and noise datasets, respectively. Each speech utterance is mixed with a random noise segment at a random Signal-to-Noise Ratio (SNR) sampled from $-20$ to $20$ dB (in 1 dB increments), to create the validation set. Clean speech data for testing are taken from the Librispeech test-clean-100 corpus. $200$ clean speech utterances are randomly picked (without replacement). $200$ noise recordings were randomly selected from the three AV datasets. Each clean speech recording is degraded by a random section from the noise recording at five SNR levels, i.e., -5 dB, 0 dB, 5 dB, 10 dB, and 15 dB. This produces 1 00010001\,0001 000 noisy AV clips for evaluation. All audio recordings are resampled at a rate of 16 kHz. The STFT spectrum is calculated using a Hanning window of 32 ms (512 samples) with a hop length of 16 ms (256 samples), resulting in a $257$-point magnitude spectrum as the acoustic input to the network. For video processing, we follow CAV-MAE [58] and employ a frame aggregation strategy to reduce computational overhead and fit our resources. In particular, we uniformly sample 10 RGB frames from each 10-second the video sequence (i.e., 1 FPS). At the training time, one RGB frame is randomly selected as the input. At the inference time, the visual embeddings of all sampled frames are averaged to generate the final scene visual embedding. Frame aggregation balances computational efficiency with the ability to capture contextual visual information, ensuring meaningful complementary cues for the audio input. In terms of time dimension, $T_{y}$ denotes the maximum time frame within a batch, while $T_{a}$ and $T_{v}$ are fixed at 128 and 49, respectively. The model dimension, $d_{model}$, is set to 256. For Mamba layers, we employ the default hyper-parameters [44]: the state dimension $d_{state}=16$, convolution dimension $d_{conv}=4$, and the expansion factor of 2.
IV-A4 Training Strategy
@@
IV-A5 Evaluation Metrics
We employ the Perceptual Evaluation of Speech Quality (PESQ[68] and the Short-Time Objective Intelligibility (STOI[69] as the evaluation metrics. Specifically, PESQ measures human-rated speech quality by predicting the Mean Opinion Score (MOS) for objective speech perception, ranging from -0.5 to 4.5, with higher values indicating better quality. STOI assesses speech intelligibility based on spectral correlation, typically ranging from 0 to 1, where higher scores reflect greater intelligibility.
@@ IV-A5 Evaluation Metrics
IV-B Experimental Results
Table I showcases the PESQ and STOI scores of the models across five SNR levels, with the best results are highlighted in bold. The evaluation results demonstrate that our VC-S2E model consistently achieves superior performance in all three datasets. In the MUSIC dataset, the unprocessed data have an average PESQ and STOI of 1.47 and 83.65%, respectively. By applying conformer-based MP-SENet, the resulting enhanced speech has improved PESQ and STOI to 2.54 and 91.65%, respectively. The AViTAR method, which employs a Conformer-based cross-attention AV framework, achieving a PESQ of 2.73 and a STOI of 92.81%. In contrast, our VC-S2E proposal shows superior performance by incorporating the BiMamba architecture with an average PESQ of 3.02 and STOI of 94.11%. It can also be seen that by integrating the spectrum, visual, and scenario-aware audio embedding, VC-S2E exceeds ExtBiMamba with an average PESQ and STOI of 2.77 and 93.09%, respectively. Moreover, the AVDCNN method leverages a dual-stream CNN with late audio-visual fusion, resulting in an average PESQ of 2.23 and STOI of 89.83%, respectively. MLDM, which utilizes a multi-level distortion measure objective, achieves an average PESQ of 2.51 and STOI of 91.22%. Meanwhile, ALTKD employs a U-Net-based teacher-student framework that integrates ultrasound tongue images and lip videos; however, it attains only an average PESQ of 2.05 and STOI of 87.83%. Notably, the DNE method, which exploits VAD to extract noise features for robust SE, records an average PESQ of 2.93 and STOI of 93.41%. Despite the competitive performance of DNE at specific SNR levels, our proposal still outperforms all these approaches, achieving an average PESQ of 3.02 and STOI of 94.11%. Furthermore, the same observations can be seen on the AVSpeech and AudioSet datasets as well, where our proposed VC-S2E achieves superior performance with the PESQ and STOI of 2.76 and 92.70% for AVSpeech, 2.59 and 92.25% for AudioSet, respectively.
Figure 4 illustrates the PESQ and STOI of our proposed VC-S2E method and its counterpart without visual and scenario-aware embeddings where AudioSet is used as the noise dataset. In both figures, VC-S2E consistently outperforms the other across all categories. For example, the PESQ of ‘Fire’ improves from 2.50 to 2.62 while for category ‘Child speech’, the score improves from 2.58 to 2.73. Similarly, the STOI scores (Figure 4(b)) also show significant improvements. For ‘Fire’, the STOI score increases from 91.61% to 92.57% while for ‘Child speech’, it improves from 92.53% to 93.27%. These results clearly reveal that incorporating visual and audio scenario-aware embeddings leads to substantial improvements in both perceptual quality and intelligibility across each individual audio category. Furthermore, we also test the model performance on five new categories (i.e., dog, insect, chewing, car and doorbell) where the results indicated in Figure 6 show the generalization of our proposed model.
((a)) PESQ scores
((b)) STOI scores
Figure 6: The (a) PESQ and (b) STOI for scenes unseen at the training time.